The mass of a neutron is 1.67×10-24 g. How many neutrons would be needed to have the same mass as a salt crystal with a mass of 2.50×10-2 g?


.............................neutrons


1
Expert's answer
2021-01-27T04:40:25-0500

The number of neutrons required to have the mass of 2.50×10-2 g equals:

n(neutrons) = (mass of neutrons) / (mass of one neutron) = 2.50 × 10-2 g / 1.67 × 10-24 g = 1.50 × 1022


Answer: 1.50 × 1022 neutrons